OpenStudy (meilendurcer):

Hallie is trying to win the grand prize on a game show. Should she try her luck by spinning a wheel with 6 equal sections labeled from 1 to 6 and hope she gets a 5, or should she roll two number cubes and hope she gets the same number on both cubes? Explain.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (meilendurcer):

what

OpenStudy (skullpatrol):

Do you have any ideas on how to solve this?

OpenStudy (meilendurcer):

no

OpenStudy (skullpatrol):

The spinning number wheel seems simple enough to find the chance that a "5" will show up, right?

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

Should she try her luck by spinning a wheel with 6 equal sections labeled from 1 to 6 and hope she gets a 5

OpenStudy (anonymous):

@jabez177 can you explain it better to me? :)

jabez177 (jabez177):

The spinner is a better application.

OpenStudy (anonymous):

i agree but how?

jabez177 (jabez177):

Because at least this one spins and dices you are never going to get the same number on two different dices. It's unlikely in the terms of probability.

OpenStudy (anonymous):

hmm tell me if this is right: soo... 6/36 for the two dices 1/6 for the spinner 6/36 = 1/6 so the spinner is better? how since they are they same?

jabez177 (jabez177):

Hold on. I'm getting mixed up. @cookiemonster2412 @dan815

OpenStudy (anonymous):

ughhh i hate probability... x|

jabez177 (jabez177):

I know! @Awolflover1

OpenStudy (cookiemonster2412):

I agree with @jabez177

OpenStudy (anonymous):

maybe... it is easier to land on 5 on a spinner than rolling two dices and landing on the same number

OpenStudy (anonymous):

so... it is easier to land on 5 on a spinner than rolling two dices and landing on the same number because 6/36 > 1/6 (the only part im confused about is the 6/36 simplified is 1/6)

OpenStudy (anonymous):

i think they have the same probability cuz 1/6 = 1/6 so either would be a good choice.